(1 point) Use Laplace transforms to solve the integral equation y(t) – v yết – U) do = 4. The first step is to apply the LaplShow work please
Answer #1

x {suces] - cflet e syry (tv)diue} = {{4} = 4 Y() - 2{est * n(t)} = 24 :. Y(4) - {€31.864) - 1. YLA) – So.VS) = Y(2) {1- 1 3